Could someone help me with this word problem.I do not have a clue on how to do this. A 150-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 30 ft longer that the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.
Answer by mananth(16946)

the diagonal brace =150 feet
the vertical height on the side of bridge = x
the horizontal height = x +30
..
they form a right triangle.
use pythagoras formula
x^2+(x+30)^2=150^2
x^2+x^2+60x+900=150^2
2x^2+60x+900=22500
2x^2+60x-21600
/2
x^2+30x-10800
x^2+120x-90x-10800=0
x(x+120)-90(x+120)=0
(x-90)(x+120)=0
x= 90 vertical distance
horizontal = 90+30 =120 feet
...
m.ananth@hotmail.ca